# Learning to Visually Connect Actions and their Effects

### v1

@@ 6.1 Cross Sample Action Selection
Ours SSL CATE Action Specification 79.36
Table 4: Performance evaluation on action quality assessment task on MTL-AQA dataset.
6.3 SSL Action Specification
In this case, we train using self-supervision grounded in connecting actions and their effects, as presented in section 5. Following the literature [34, 31], we carry out SSL training on the training set of the MTL-AQA dataset [30]. After SSL training, we do not perform any fine-tuning; we simply use the backbone as a feature extractor. Features are extracted from 16 uniformly sampled frames, and then a linear regressor is applied on top of them. We compare SSL CATE-action specification with other self-supervised models, and the results are shown in Table 4. We observe that SSL-CATE action specification self-supervision outperforms other state-of-the-art methods, indicating that CATE-Action Specification does help in learning features suitable for fine-grained action analysis tasks such as action quality assessment.
Figure 8: Pose retrieval results. Top-3 retrievals containing diver using our model vs [44]. In this experiment, what we are looking for is if the retrieved results have the athlete in similar pose and orientation as the query pose (left most column). For explanation, we have drawn red bounding box around wrong retrievals. As highlighted with white-colored annotations in zoomed in view, in the wrong retrieval the athlete is in somersault tuck pose, while the query has the athlete in pike position. In the correct retrieval as done by our SSL-CATE-Action Specification, the retrieval results have the athlete in the pike position.
SSL-CATE action specification trained representations learn to encode human pose.
In this experiment, we further probe into what the SSL-CATE Action Specification network might be paying attention to. Using the trained model from the above experiment, features are extracted from diving video clip frames. Then, we carried out a pose retrieval experiment—given a frame containing a diver in a certain pose (query pose), the objective is to retrieve divers in a similar pose as the query pose. We further compare retrieval done using SSL-CATE trained representations against an off-the-shelf dedicated pose estimation model [44]. We find that CATE representations are better able to retrieve poses than a dedicated pose estimation model (Figure 8). We hypothesize that this might be because the off-the-shelf pose estimator might not be able to handle motion blurring and divers in convoluted poses, while CATE-based self-supervision allows learning pose-sensitive robust representations directly from challenging videos.
@@ 7 Conclusion
In this work, we address the underexplored area of visually connecting actions and their effects in video understanding. By introducing tasks like action selection and action specification, we shed light on the challenges and opportunities in understanding causal mechanisms for machines. Despite efforts to enhance state-of-the-art models, a significant performance gap between human and machine capabilities persists in this task. The study lays a solid foundation for future research, emphasizing the flexibility and versatility of connecting actions and effects. We hope that these insights will inspire the development of advanced formulations and models, paving the way for improved performance and broader applications in the field of video understanding.

### v3

@@ 6.3 SSL Effect-Affinity Assessment for AQA
Ours SSL CATE Effect-Affinity Assessment 79.36
Table 3: Performance evaluation on action quality assessment task on MTL-AQA dataset. Performance metric: Spearman’s rank correlation coefficient (%).
Figure 7: Pose retrieval results. Top-3 retrievals containing diver using our model vs [52]. In this experiment, what we are looking for is if the retrieved results have the athlete in similar pose and orientation as the query frame (left most column). For explanation, we have drawn red bounding box around wrong retrievals. As highlighted with white-colored annotations in zoomed in view, in the wrong retrieval the athlete is in somersault tuck pose, while the query has the athlete in pike position. In the correct retrieval as done by our SSL-CATE-Effect-Affinity Assessment, the retrieval results have the athlete in the pike position.
SSL-CATE Effect-Affinity Assessment trained representations learn to encode human pose. In this experiment, we further investigate what the SSL-CATE Effect-Affinity Assessment network might be focusing on. Using the trained model from the previous experiment, features are extracted from frames of diving video clips. We then conducted a pose retrieval experiment. Given a frame containing a diver in a certain pose (query pose), the goal is to retrieve frames of divers in similar poses and orientations as the query pose based on feature similarity. We compared the retrieval performance of SSL-CATE trained representations against an off-the-shelf dedicated pose estimation model [52]. Our findings indicate that CATE representations are more effective at retrieving similar poses than the dedicated pose estimation model (Figure 7). This better performance suggests that SSL-CATE Effect-Affinity Assessment representations encode fine-grained details like the pose of the person. We hypothesize that this may be because the off-the-shelf pose estimator struggles with motion blur and convoluted poses, while the CATE-based self-supervision learns pose-sensitive and robust representations directly from challenging videos.
@@ 7 Conclusion
In this work, we explore the concept of connecting actions and their effects (CATE), which has applications in learning from demonstration and task planning. We identify and study two key aspects: Action Selection (AS) and Effect Affinity Assessment (EAA), formalize them, introduce datasets, and develop baseline solutions. While our baselines improve over existing models, a significant performance gap remains between humans and machines on AS. Our experiments show that in solving AS & EAA, models learn intuitive properties like object tracking and pose encoding without explicit supervision. We demonstrate that CATE can be an effective self-supervised task for learning video representations from unlabeled videos. CATE is a fundamental cognitive process with potential applications in learning from demonstrations, human-machine interactions, and creating adaptive AI tools. Our work lays the foundation for future innovations, inspiring advanced models with broader applications and improved performance.
@@ Acknowledgements.
Appendix
8 Dataset samples with and without object detections
For easier viewing, we have provided them in the accompanying Video.
9 Further details on Baseline Methods for Action Selection
In this section, we provide further details on the various baseline approaches that we design and evaluate. Actions are represented as video clips, while states by very short clips. Unless specified otherwise, we use a backbone model ($\mathcal{E}$) pretrained on a large-scale action recognition dataset to extract features for actions and states. Subsequently, various modules mentioned in the following operate on these features.
9.1 Naive model
Approach is illustrated in Figure 8. As a simple baseline, in naive matching the average of the initial and desired or the final state features are matched with the action features using cosine similarity. During inference, the action option with the highest similarity score is selected as the correct answer. Note that, this model involves no training; pretrained action recognition model is used as the state and action feature extractor.
Figure 8: Naive model. This model involves no training, but only inference. $\bigoplus$ represents averaging operation.
9.2 Treating Actions as Transformations
[…]
The approach (in training mode) is shown in Figure 9. Actions can be viewed as transformations [50] which when applied to the initial state yields the desired or the final state. While the original approach uses fixed transformations [50], our adaptation facilitates conditional transformations computed on the fly. We develop a model in which a transformation vector is computed from an action clip using a dense network. This transformation vector ($A$) when multiplied with the initial state vector ($I$) yields the final state vector ($F$). During training, we use contrastive learning [4] to train the parameters of action encoder: Cosine similarity between predicted and actual desired state representations is maximized for the correct action and minimized for the incorrect action. During inference, the action option predicting the final state with the highest similarity to the given final state feature is selected as the correct answer. We term these models as Actions as transformations (AT).
Figure 10: MoRISA model.
9.3 Modeling rich interactions between initial state & action (MoRISA)
Approach (in training mode) is shown in Figure 10. Here a bilinear model [26] is leveraged to capture rich interactions between the initial state and the action to generate the final state in feature space. Essentially, initial state ($I$) and action representations ($A$) are fed into a bilinear model, which outputs a final state representation ($F$). The parameters of the model are trained using contrastive learning as before and the inference is done similarly to AT model above.
(a) LinSAES Concept.
(b) LinSAES model.
Figure 11: LinSAES concept (a) and model (b).
9.4 LinSAES: Learning Linear State-Action Embedding Space
Concept of this approach is shown in 11(a); and the framework (in training mode) is shown in 11(b). In this framework, actions and initial states undergo transformation into a linear embedding state-action representation space through a Transformer encoder (we term it as AEXformer in 11(b)). The hypothesis suggests that within this linear, disentangled space, adding an action vector ($A$) to the initial state representation ($I$) enables a transition to the final state in the representation space ($F$) (illustrated in conceptual diagram in 11(a)). The Transformer encoder processes a sequence of initial state and action representation vectors, and the resulting output is decoded by a linear decoder ($\mathcal{D}$) to produce the final state feature vector. During training, all model parameters are optimized end-to-end with the objective of aligning the resultant final state representation with the ground truth final state representation. Optimization and inference process remains the same as AT model. We call this model LinSAES.
Figure 12: Connecting via Swapping. Here, we have not shown using of incorrect action for better explanation and avoiding confusion. However, in practice, we leverage incorrect actions.
9.5 Connecting via Swapping
We will start by presenting method overview, followed by details. Approach (in training mode) is illustrated in Figure 12.
Method overview:
all model parameters (encoders ($\alpha,\beta$) and decoders ($\mathcal{D}$)) are trained end-to-end by enforcing the following. 1) Transformation Encoding. The initial and final states are transformed into an initial state and a corresponding state-transformation code through an encoder. This transformation code captures the changes or effects between the initial and final states. 2) Reconstruction using Decoder. Given the initial state and the state-transformation code, a decoder reconstructs the final state. This decoder essentially learns to generate the final state based on the initial state and the transformation code. 3) Obtaining Transformation Code from Cross-sample Actions. Additionally, a state-transformation code can be obtained from the action (but from another sample) occurring between the initial and final states. This is done through another encoder that distills the transformation code from the action. Note that, during training the final state is reconstructed using both transformation codes—obtained from states and action. By obtaining equivalent state-transformation codes both from the states directly and from the actions, the model learns to connect actions to their effects. The decoder then utilizes this transformation code to reconstruct the final state from the initial state, effectively capturing the impact of actions on the state transition in the video data.
Method details:
To explicitly isolate a $n$-dimensional “transformation “code”, we pass the initial and final states (each represented by $m$-dimensional features) through an encoder $\alpha$, which outputs a $(m+n)$-dimensional vector. We enforce the condition that the first $m$ elements of this vector represent the initial state and the remaining $n$ elements represent the transformation “code”. Now, using another encoder ($\beta$), we distill a $n$-dimensional transformation code from an action option. We swap the transformation code obtained using states with that obtained from the action option. A decoder ($\mathcal{D}$) then reconstructs the final state from the extracted initial state and swaps the transformation code. All encoders and decoders are trained by enforcing that the extracted initial state be similar to the ground truth initial state and the reconstructed final state be similar to the ground truth final state in case the transformation code was coming from the correct action option (minimize the similarity if the transformation code was from incorrect action option). This approach is inspired by [34]. However, while their approach disentangles human pose and appearance, our approach extends it to disentangle state-transformation for linking actions and their effects.
9.6 CLIP
For this baseline, we adopt the previously discussed naive approach subsection 9.1. In this case, we use the CLIP [38] representations for states and actions. Further, to obtain a clip-level representation for action-clips, we average the frame-level representations. Inference process remains as the Naive approach.
9.7 VideoChat2
VideoChat2 [24] is powerful video understanding foundational model. It is trained on 20 challenging video understanding simultaneously. We evaluated various ways to adopt VideoChat2 for our problem with the objective of maximizing the performance on Action Selection. We obtained best results with the following approach. First, we provide the initial and the final states to the model alongwith the four action classes in language format, ask to identify the action taking place. Second, we ask the model to identify the action class (from the four choices in language format—these are nothing but the class names of the four action options) taking place in the correct action-clip. Then, we match the classes obtained from states and action-clips. If there is a match and the predicted class is correct, we consider that the model got the answer, else, we consider the answer of the model as wrong.
9.8 VideoMAE
Mask autoencoding (MAE) might seem similar to Action Selection, but there is a fundamental difference between them. MAE randomly masks patches without any explicit attention to initial and final states, whereas Action Selection particularly involves connecting actions and desired state changes. Nonetheless, VideoMAE has achieved strong results in video understanding. Therefore, for completeness, we consider VideoMAE as one of the baselines. By design, VideoMAE cannot be directly used for our task, so we adopt the previously discussed Naive approach, but using VideoMAE as the backbone for Action Selection task.
10 Introduction to Action Quality Assessment (AQA) and how it is connected to Effect-Affinity Assessment
Definition of AQA. AQA is a fine-grained action analysis task, where the models try to assess how well an action was performed. Judging during Olympics diving can be a classic example of AQA. AQA requires paying attention to very fine-grained details of action. For example, it involves analyzing how tight was the athlete’s form during somersaults, how close the athlete’s feet were during twisting motion in air, was the athlete under-/over-rotated during entry into the water, how high the athlete jumped during take-off, etc. Taking these into consideration, a score is given out to indicate how good the athletes’ performance was or whether the rules of the competition were followed.
Effect-Affinity Assessment and AQA. In our formulation, Effect-Affinity Assessment involves discerning between very nearby effects, and determining how far is an effect-frame from the action applied on an initial state. To solve this task, the model learns to pay attention to fine-grained details like what maneuvers the athlete applied/executed, and how that would result in what series of poses, or how the athlete’s position would evolve as a result of their maneuvers. These details have an overlap with the elements of interest in AQA. Therefore, we hypothesize that representations learnt in solving our SSL CATE Effect-Affinity Assessment should transfer well to AQA task.
11 Action-Effect Joint Attention Visualization
We use the best performing Analogical-Reasoning model for attention visualization, employing a modified GradCAM [41] to generate joint attention heatmaps over states & actions. Specifically, we backpropagate the dot-product of state-change vector & the action vector through the initial, final, action branches into visual input space.
12 Further Action Selection performance analysis
We have also provided the classwise accuracies from two of the best performing models: 1) Analogical Reasoning model; 2) LinSAES (learning linear state-action embedding space) in the Figure 13. In both cases, we observed a similar trend that the action classes on which the models were most accurately connected actions and effects were simple and single-step movements. Action classes where the models struggled the most were composite and complex involving multiple steps or sub-actions. These composite actions may require the model to understand not only individual actions but also their sequential dependencies and temporal relationships. The increased complexity can make it more challenging for the model to accurately predict the correct sequence of actions to achieve the desired outcome. The easiest classes to connect were derived from SSv2 dataset [12] representing common tasks or actions one might encounter in daily life or simple mechanical tasks. These actions focus on the manipulation and movement of objects in various ways, often on flat surfaces. Most difficult to connect actions were from COIN dataset [45], often associated with particular activities such as cooking, construction, or assembly. These actions involve more specific and varied tasks such as cutting, filling, inserting, and cooking, indicating a wider scope of activities.
Figure 13: Classwise accuracies from Analogical Reasoning (AR) LinSAES (AEX) models. Easiest and most difficult classes are listed below.